OpenStudy (anonymous):

x is less than or equal to -3 whAT are the numbers?

OpenStudy (math_bender_15):

\[-3\le x\]

OpenStudy (anonymous):

\[x \le-3\]

OpenStudy (anonymous):

What is the question? You seem to know that it is every number less than -3 including -3. So -3, -4, -5 ,-6 etc.

OpenStudy (anonymous):

im confused in graphing, im sorry, hehe, thank you, is it the same with the \[3\le x\]

OpenStudy (anonymous):

Sorry I don't really understand. Is that a separate question?

OpenStudy (anonymous):

yeah hehe,,